OpenStudy (anonymous):

Determine the zeros of f(x) = x^3 - 3x^2 - 16x + 48 show work please

OpenStudy (apoorvk):

\[=x^2(x-3) - 16(x-3)\] Can you proceed now?

OpenStudy (anonymous):

what is that?

OpenStudy (anonymous):

could you plug it in please

OpenStudy (apoorvk):

Just proceed like you do it for a normal binomial-binomial factorisation.

OpenStudy (anonymous):

factoring by grouping take largest term you can factor from left side (x^2) and largest term you can factor from right side (16), and factor them out.

OpenStudy (anonymous):

zack, look at what @apoorvk 's last line... you have common factors of (x-3)... factor those out...

OpenStudy (anonymous):

what now?

OpenStudy (anonymous):

=x2(x−3)−16(x−3) what after this? is the answer -4?

OpenStudy (anonymous):

You see, you factor by grouping first: f(x) = x^3 - 3x^2 - 16x + 48 f(x) = x^2(x - 3) - 16(x - 3) f(x) = (x^2 - 16)(x - 3) Now, there is a difference in squares pattern in there. f(x) = (x^2 - 16)(x - 3) f(x) = (x + 4)(x - 4)(x - 3) Set it equal to zero. 0 = (x + 4)(x - 4)(x - 3) Now solve for x in each of these binomials by setting it equal to zero. x + 4 = 0 x = -4 x - 4 = 0 x = 4 x - 3 = 0 x = 3 Therefore, the roots are x = -4, 4, 3.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome. Do you understand it?